# Measuring What Matters!! Assessing Therapeutic Principles in Mental-Health Conversation

### 07 Apr 2026 v1

@@ Abstract
Abdullah Mazhar1, Het Riteshkumar Shah1, Aseem Srivastava2,thanks: Work done during his tenure at FLaME Lab, IIIT Delhi., Smriti Joshi3, Md. Shad Akhtar1 1IIIT Delhi, India, 2MBZUAI, UAE, 3Wysa Correspondence: abdullahm@iiitd.ac.in
@@ 2 Related Work
A substantial body of work focuses on using AI to detect mental health issues from textual data. Studies have used social media and clinical transcripts to identify depressive or anxiety-related symptoms using supervised learning and lexicon-based techniques Yadav et al. (2020); Cha et al. (2022); William and Suhartono (2021). The DAIC corpus was also introduced, incorporating verbal and non-verbal signals to support multimodal depression detection. While impactful for early screening, these systems are not designed for interactive therapeutic use and often lack contextual reasoning or dialogic modeling Gratch et al. (2014).
@@
Conversational agents are increasingly explored for supportive roles in mental health contexts. The coherence of LLM responses in therapeutic dialogues has been examined Song et al. (2024), and ChatGPT responses have been evaluated against therapeutic principles Hatch et al. (2025). Other efforts focus on counseling summarization Srivastava et al. (2022); Chen and Verma (2006); Srivastava et al. (2024), dialogue understanding Malhotra et al. (2022); Saha et al. (2021), and response generation Srivastava et al. (2023); Lippe et al. (2020). End-to-end systems such as CareBot Crasto et al. (2021) aim to offer structured support. However, most evaluations emphasize fluency, empathy, or informativeness, relying on general-purpose metrics or subjective judgments rather than clinically grounded assessment.
@@
Empathy is widely recognized as central to therapeutic communication. Prior work has explored computational empathy in online forums Sharma et al. (2021), peer-support conversations Srivastava et al. (2025), and dialogue systems Cai et al. (2024). Approaches range from affective classification to empathy-driven generation using transformer models Goel et al. (2021); Zaranis et al. (2021). However, empathy alone does not guarantee therapeutic safety or professional adherence, as it may overlook essential principles such as respect for autonomy or situational appropriateness.
@@ 3.3 Data Analysis
FAITH-M comprises 10,172 utterances from 167 dialogues and is annotated with six ordinal labels corresponding to core therapeutic principles. The dataset is split session-wise into train (6,902 utterances from 116 sessions), validation (949 utterances from 17 sessions), and test (2,321 utterances from 34 sessions) sets in a 70:10:20 ratio. The detailed distribution for each therapeutic dimension is presented in Table 5 (c.f Appendix).
@@ 4.2 KD-CoT
We embed each patient-therapist pair using a Sentence Transformer Reimers and Gurevych (2019). For a given test instance, we compute cosine similarity between it’s embedding and those of the reference sets. We then retrieve the top two most semantically similar reference pairs per dimension and polarity. The aim is to keep the selected examples both contextually relevant and therapeutically aligned. The retrieved exemplars from the previous step are passed to GPT-4o along with a structured prompt [^4] Refer to Appendix to view prompts (cf Figure 7).. GPT-4o is used to generate dimension-specific Chain-of-Thought explanations grounded in therapeutic norms. Figure 5 (c.f Appendix)
@@ 4.3 Ordinal Classification Block
We integrate the context-aware representation $R_{\text{ctx}}$, the distilled knowledge representation $R_{\text{KD}}$, and the utterance embedding $r_{t}$ using a cross-attention-based fusion mechanism. Specifically, $r_{t}$ is used as the query, while $R_{\text{ctx}}$ and $R_{\text{KD}}$ serve as key-value pairs, enabling the model to attend over contextual and knowledge signals before ordinal classification. We then pass the resulting fused representation to an ordinal classification head to predict alignment scores across five levels. To jointly model ordinal structure and categorical accuracy, we use a hybrid loss:
@@ 5 Experiments and Results
We benchmark CARE against a diverse set of 15 baselines, comprising 4 prompt-based and 11 trainable models, evaluated under a unified experimental setup. The baseline methods can be viewed in Table 2, and we include more details pertaining to the baseline checkpoints in Section C (c.f. Appendix). All baselines use the same local context window ($k=2$) and, where applicable, are trained with the same ordinal-aware loss function (Equation 1).
Performance Comparison.
@@
Ablation Study.
We conduct an ablation study to assess the contribution of core CARE components, including local contextual encoding and the KD-CoT module. Here, GPT-4o and MentalLLaMA are used as reasoning teachers to generate Chain-of-Thought rationales within the KD-CoT framework, while the final predictions are produced by the student model. Table 3 reports performance as we progressively add contextual grounding, label-exclusive exemplar retrieval, and contrastive KD-CoT reasoning, showing consistent improvements with each component. The largest gains arise from combining context with polarity-aware exemplar reasoning, indicating that CARE’s effectiveness stems from structured reasoning rather than dialogue context or prompt supervision alone. Figure 6 visualizes these trends, where moderate context sizes (k=2k=233) yield peak performance and larger windows show diminishing returns. Additional ablation analyses and implementation details are provided in (Appendix  G).
6 Analysis
Generalizability.
In addition to the CARE’s benchmarking on FAITH-M (c.f Table 2), we extended the CARE’s inference on two cross-dataset evaluations: PTSD Suhas et al. (2025) and Cheeseburger  [^5] https://cheeseburgertherapy.org/. Table 9 (c.f Appendix) shows detailed results and highlights that the advantage persists under cross-dataset evaluation. On the PTSD dataset, which contains trauma-focused Prolonged Exposure therapy dialogues, and the CBT-based Cheeseburger dataset, CARE consistently outperforms prompt-based and fine-tuned LLM baselines. Although absolute performance decreases relative to in-domain evaluation, as expected with considerable domain-shift, the relative improvements remain stable, suggesting that CARE generalizes beyond the distribution on which it was developed.
Quantitative Analysis.
We analyze the confusion matrices to understand error patterns across therapeutic dimensions using the updated model configuration (Figure 4). For the Reflecting Feelings dimension, the model shows reliable recognition of extreme categories, with accuracies of 66% for Strong Negative and 49% for Strong Positive. As expected in ordinal classification, most errors occur between adjacent labels, particularly between Mild Positive and Strong Positive (30%), reflecting the difficulty of distinguishing subtle emotional affirmation from strong empathic reflection. Similar trends are observed across the remaining therapeutic principles (Appendix, Figure 9). The model consistently maintains strong diagonal mass for extreme labels, achieving approximately 65–70% accuracy for Strong Negative and Mild Negative classes across dimensions such as Non-Judgmental Language, Active Listening, and Situational Appropriateness. In contrast, mid-spectrum categories remain challenging, with neutral labels often confused with adjacent mild positive or mild negative classes (15-20% misclassification, depending on the dimension). Overall, these confusion patterns indicate that the model captures ordinal structure effectively, with errors largely constrained to neighboring categories rather than cross-polarity confusions. This behavior suggests sensitivity to graded therapeutic alignment while highlighting the inherent ambiguity of mid-spectrum judgments.
@@ 6 Analysis
Table 4 presents cases where CARE diverges from expert annotations. Errors typically arise from misinterpreting subtle therapeutic cues, such as emotionally grounded paraphrasing, which the model sometimes labels as overly positive or neutral, consistent with patterns observed in earlier analyses (c.f. Section 6). Across $6$ utterances, the model deviated from gold labels in $7$ out of $18$ dimension-level predictions, usually by a margin of $\pm 1$. For example, the utterance “So you consider yourself Catholic. But you feel like there’s a lot of rules.” was annotated as neutral for non-judgmental acceptance, while the model predicted mild-positive, reflecting a tendency to overestimate surface-level reflectiveness as affirming. Errors were most frequent for situational appropriateness and reflective feelings (3/6 cases each), whereas non-judgmental acceptance and respect for autonomy showed higher agreement (4/6 correct each). The model occasionally carries forward prior sentiment across topic shifts, leading to misalignment in situational appropriateness and highlighting challenges in modeling mid-range ordinal labels.
@@ Discussion
We analyze GPT-4o under zero-shot and few-shot prompting to examine the limitations of prompt-based therapeutic evaluation. While few-shot prompting yields a modest gain in average accuracy, it leads to a substantial drop in weighted F1, indicating poorer calibration across ordinal categories. This suggests that in-context examples improve format adherence without enhancing sensitivity to ordinal severity. Per-dimension results (Appendix Table 10) show stronger performance on Warmth and Encouragement and Active Listening, where alignment is often conveyed through surface-level cues, and weaker performance on Non-Judgmental Language and Respect for Autonomy, which require reasoning about implicit judgment and relational intent, as evident under zero-shot prompting (Figure 10), GPT-4o frequently collapses negative and neutral instances into the Neutral class. Few-shot prompting (Figure 11) shifts predictions toward mild and strong positive labels without recovering negative or neutral categories. Figure 8 (c.f. Appendix) illustrates GPT-4o’s tendency to replicate exemplar patterns rather than adapt to the target context, highlighting structural limits of prompt-only evaluation for fine-grained therapeutic assessment.
@@ 9 Acknowledgment
Appendix A Data Distribution & Expert Agreement
Therapeutic Dimension SN MN Neu MP SP
Non-Judgmental Acceptance 107 413 624 1133 771
Warmth and Encouragement 42 217 733 1134 921
Respect for Autonomy 112 208 502 1401 825
Active Listening 54 89 546 1458 900
Reflective Feelings 84 155 116 1056 589
Situational Appropriateness 79 151 290 1279 1248
NJL Warmth Autonomy Listening Reflect SA
81.60% 84.00% 70.30% 70.10% 66.70% 69.20%
Table 6: Agreement score: Expert vs CARE.
Few-Shot vs Zero-Shot Prompt Design for GPT-4o Evaluation Few-Shot Prompt: Instruction: You are a clinical NLP evaluator. Score a therapist response from -2 to +2 across six therapeutic principles. You will be shown few-shot examples and then asked to score a new interaction. ### Example Patient: <previous utterance> Therapist: <therapist response> Explanations: - Non-Judgmental Language: ... - Warmth and Encouragement: ... - Respect for Autonomy: ... - Active Listening: ... - Reflecting Feelings: ... - Situational Appropriateness: ... Scores: NJ: X, W: X, RA: X, AL: X, RF: X, SA: X --- Repeat for multiple examples. ### Now evaluate this response: Patient: <current patient utterance> Therapist: <current therapist response> Explanations: ... Return scores in same format. Zero-Shot Prompt: Instruction: You are a mental health evaluation assistant trained in psychotherapy communication techniques. Please assess the quality of the therapist’s response using the following six therapeutic dimensions. Use the preceding patient’s statement to help you understand the context and appropriateness of the therapist’s message. --- Patient said: "<patient utterance>" Therapist responded: "<therapist utterance>" --- Rate the therapist’s response from -2 to +2 for each dimension: - -2 = strongly violates the principle - 0 = neutral or unclear - +2 = strongly demonstrates the principle Output Format (verbatim): Non-Judgmental: X, Warmth: X, Respect: X, Active Listening: X, Reflecting Feelings: X, Situational Appropriateness: X
Appendix B Evaluation Metrics and Loss.
We evaluate model performance using standard metrics: precision (Pre), recall (Re), and weighted F1-score ($F1_{w}$), which is computed per therapeutic principle across five ordinal categories. For fairness and consistency across experiments, all baselines as well as the proposed model are trained using the same ordinal-aware loss function (c.f Equation 1) that weighs both cross-entropy and mean squared error losses.
Appendix C Baseline Methods
@@
(a) Zero-shot LLMs: GPT-4o Islam and Moushi (2024) and MentalLLaMA Yang et al. (2024) are evaluated without task-specific fine-tuning to gauge the upper bound of out-of-the-box performance. (b) Few-shot LLMs: We further extend this by evaluating GPT-4o under few-shot prompting 10, using carefully constructed in-context examples (Appendix 5) to examine whether performance improves with light in-context supervision. A detailed analysis of both zero-shot and few-shot performance including prompt formats, per-dimension scores, and failure analysis is presented in (Discussion 6). (c) Transformer encoders: MentalBERT Ji et al. (2021), ALBERT Lan et al. (2019), DeBERTa He et al. (2020), and RoBERTa Liu et al. (2019) test how well discriminative, parameter-efficient encoders capture therapeutic nuances when fine-tuned on our data; (d) Encoder–decoders: BART Lewis et al. (2019) and its domain-adapted variant MentalBART Yang et al. (2023) assess the utility of sequence-to-sequence pretraining for dialogue-level classification tasks; (e) Decoder-only LLMs: Qwen3 Yang et al. (2025), LLaMA 3.1/3.2 Grattafiori et al. (2024), Phi-4 Abdin et al. (2025), and Gemma Team et al. (2024) probe the limits of autoregressive generators when repurposed for structured ordinal prediction.
Appendix D Experimental Setup

### 13 Apr 2026 v2

@@ Abstract
Abdullah Mazhar, Het Riteshkumar Shah, Aseem Srivastava, Smriti Joshi, Md. Shad Akhtar IIIT Delhi, India; MBZUAI, UAE; Wysa {abdullahm, het22213, shad.akhtar}@iiitd.ac.in, smriti@touchkin.com, aseem.srivastava@mbzuai.ac.ae
@@ 2 Related Work
A substantial body of work focuses on using AI to detect mental health issues from textual data. Studies have used social media and clinical transcripts to identify depressive or anxiety-related symptoms Mazhar et al. (2025) using supervised learning and lexicon-based techniques Yadav et al. (2020); Cha et al. (2022); William and Suhartono (2021). The DAIC corpus was also introduced, incorporating verbal and non-verbal signals to support multimodal depression detection. While impactful for early screening, these systems are not designed for interactive therapeutic use and often lack contextual reasoning or dialogic modeling Gratch et al. (2014).
@@
Conversational agents are increasingly explored for supportive roles in mental health contexts Srivastava et al. (2025c). The coherence of LLM responses in therapeutic dialogues has been examined Song et al. (2024), and ChatGPT responses have been evaluated against therapeutic principles Hatch et al. (2025). Other efforts focus on counseling summarization Srivastava et al. (2022); Chen and Verma (2006); Srivastava et al. (2024), dialogue understanding Malhotra et al. (2022); Saha et al. (2021), and response generation Srivastava et al. (2023); Lippe et al. (2020). End-to-end systems such as CareBot Crasto et al. (2021) aim to offer structured support. However, most evaluations emphasize fluency, empathy, or informativeness, relying on general-purpose metrics or subjective judgments rather than clinically grounded assessment.
@@
Empathy is widely recognized as central to therapeutic communication. Prior work has explored computational empathy in online forums Sharma et al. (2021), peer-support conversations Srivastava et al. (2025a), and dialogue systems Cai et al. (2024); Srivastava et al. (2025b). Approaches range from affective classification to empathy-driven generation using transformer models Goel et al. (2021); Zaranis et al. (2021). However, empathy alone does not guarantee therapeutic safety or professional adherence, as it may overlook essential principles such as respect for autonomy or situational appropriateness.
@@ 3.3 Data Analysis
FAITH-M comprises 10,172 utterances from 167 dialogues and is annotated with six ordinal labels corresponding to core therapeutic principles. The dataset is split session-wise into train (6,902 utterances from 116 sessions), validation (949 utterances from 17 sessions), and test (2,321 utterances from 34 sessions) sets in a 70:10:20 ratio. The detailed distribution for each therapeutic dimension is presented in Table 2.
Therapeutic Dimension SN MN Neu MP SP
Non-Judgmental Acceptance 107 413 624 1133 771
Warmth and Encouragement 42 217 733 1134 921
Respect for Autonomy 112 208 502 1401 825
Active Listening 54 89 546 1458 900
Reflective Feelings 84 155 116 1056 589
Situational Appropriateness 79 151 290 1279 1248
@@ 4.2 KD-CoT
We embed each patient-therapist pair using a Sentence Transformer Reimers and Gurevych (2019). For a given test instance, we compute cosine similarity between it’s embedding and those of the reference sets. We then retrieve the top two most semantically similar reference pairs per dimension and polarity. The aim is to keep the selected examples both contextually relevant and therapeutically aligned. The retrieved exemplars from the previous step are passed to GPT-4o along with a structured prompt [^4] Refer to Appendix to view prompts (cf Figure 7).. GPT-4o is used to generate dimension-specific CoT explanations grounded in therapeutic norms Figure 6 (c.f. Appendix).
@@ 4.3 Ordinal Classification Block
We integrate the context-aware representation $R_{\text{ctx}}$, the distilled knowledge representation $R_{\text{KD}}$, and the utterance embedding $r_{t}$ using a cross-attention-based fusion mechanism. Specifically, $r_{t}$ is used as the query, while $R_{\text{ctx}}$ and $R_{\text{KD}}$ serve as key-value pairs, enabling the model to attend over contextual and knowledge signals. We then pass the resulting fused representation to an ordinal classification head to predict alignment scores across five levels. To jointly model ordinal structure and categorical accuracy, we use a hybrid loss:
@@ 5 Experiments and Results
We benchmark CARE against a diverse set of 15 baselines, comprising 4 prompt-based and 11 trainable models, evaluated under a unified experimental setup. The baseline methods can be viewed in Table 3, and we include more details pertaining to the baseline checkpoints in Appendix A. All baselines use the same local context window ($k=2$) and, where applicable, are trained with the same ordinal-aware loss function (Equation 1).
@@
Performance Comparison.
Ablation Study.
We conduct an ablation study to assess the contribution of core CARE components, including local contextual encoding and the KD-CoT module. Here, GPT-4o and MentalLLaMA are used as reasoning teachers to generate Chain-of-Thought rationales within the KD-CoT framework, while the final predictions are produced by the student model. Table 4 reports performance as we progressively add contextual grounding, label-exclusive exemplar retrieval, and contrastive KD-CoT reasoning, showing consistent improvements with each component. The largest gains arise from combining context with polarity-aware exemplar reasoning, indicating that CARE’s effectiveness stems from structured reasoning rather than dialogue context or prompt supervision alone. Figure 5 visualizes these trends, where moderate context sizes (k=2k=233) yield peak performance and larger windows show diminishing returns. Additional ablation analyses and implementation details are provided in (Appendix  C).
Model Acc. Macro Weighted
F1 Prec. Rec. F1 Prec. Rec.
Dataset: PTSD
GPT-4o 38.26 25.53 28.79 28.49 35.06 37.09 38.26
Qwen 38.00 16.80 19.10 25.30 31.10 34.90 38.00
Llama 3.2 29.80 8.80 8.60 18.90 17.50 14.50 29.80
CARE 52.34 46.80 48.80 46.50 50.90 51.30 52.40
Dataset: CheeseBurger
GPT-4o 26.33 12.62 19.61 10.14 31.50 42.22 26.33
Qwen 28.05 11.25 10.71 16.58 26.94 29.35 28.05
Llama 3.2 23.15 9.01 16.98 17.89 18.95 39.13 23.15
CARE 48.96 20.46 28.28 20.07 50.30 65.25 48.96
6 Analysis
Generalizability.
In addition to the CARE’s benchmarking on FAITH-M (c.f. Table 3), we extended the CARE’s inference on two cross-dataset evaluations: PTSD Suhas et al. (2025) and Cheeseburger [^5] https://cheeseburgertherapy.org/. Table 5 shows detailed results and highlights that the advantage persists under cross-dataset evaluation. On the PTSD dataset, which contains trauma-focused Prolonged Exposure therapy dialogues, and the CBT-based Cheeseburger dataset, CARE consistently outperforms prompt-based and fine-tuned LLM baselines. Although absolute performance decreases relative to in-domain evaluation, as expected with considerable domain-shift, the relative improvements remain stable, suggesting that CARE generalizes beyond the distribution on which it was developed.
@@ 6 Analysis
Quantitative Analysis.
We analyze the confusion matrices to understand error patterns across therapeutic dimensions using the updated model configuration in Figure 4. For the Reflecting Feelings dimension, the model shows reliable recognition of extreme categories, with accuracies of 66% for Strong Negative and 49% for Strong Positive. As expected in ordinal classification, most errors occur between adjacent labels, particularly between Mild Positive and Strong Positive (30%), reflecting the difficulty of distinguishing subtle emotional affirmation from strong empathic reflection. Similar trends are observed across the remaining therapeutic principles (Appendix, Figure 10). The model consistently maintains strong diagonal mass for extreme labels, achieving approximately 65–70% accuracy for Strong Negative and Mild Negative classes across dimensions such as Non-Judgmental Language, Active Listening, and Situational Appropriateness. In contrast, mid-spectrum categories remain challenging, with neutral labels often confused with adjacent mild positive or mild negative classes (15-20% misclassification, depending on the dimension). Overall, these confusion patterns indicate that the model captures ordinal structure effectively, with errors largely constrained to neighboring categories rather than cross-polarity confusions. This behavior suggests sensitivity to graded therapeutic alignment while highlighting the inherent ambiguity of mid-spectrum judgments.
@@
Table 6 presents cases where CARE diverges from expert annotations. Errors typically arise from misinterpreting subtle therapeutic cues, such as emotionally grounded paraphrasing, which the model sometimes labels as overly positive or neutral, consistent with patterns observed in earlier analyses. Across $6$ utterances, the model deviated from gold labels in $7$ out of $18$ dimension-level predictions, usually by a margin of $\pm 1$. For example, the utterance “So you consider yourself Catholic. But you feel like there’s a lot of rules.” was annotated as neutral for non-judgmental acceptance, while the model predicted mild-positive, reflecting a tendency to overestimate surface-level reflectiveness as affirming. Errors were most frequent for situational appropriateness and reflective feelings (3/6 cases each), whereas non-judgmental acceptance and respect for autonomy showed higher agreement (4/6 correct each). The model occasionally carries forward prior sentiment across topic shifts, leading to misalignment in situational appropriateness and highlighting challenges in modeling mid-range ordinal labels.
@@
NJL Warmth Autonomy Listening Reflect SA
81.60% 84.00% 70.30% 70.10% 66.70% 69.20%
Table 7: Agreement score: Expert vs CARE.
@@ Discussion
We analyze GPT-4o under zero-shot and few-shot prompting to examine the limitations of prompt-based therapeutic evaluation. While few-shot prompting yields a modest gain in average accuracy, it leads to a substantial drop in weighted F1, indicating poorer calibration across ordinal categories. This suggests that in-context examples improve format adherence without enhancing sensitivity to ordinal severity. Per-dimension results in Table 8 show stronger performance on Warmth and Encouragement and Active Listening, where alignment is often conveyed through surface-level cues, and weaker performance on Non-Judgmental Language and Respect for Autonomy, which require reasoning about implicit judgment and relational intent, as evident under zero-shot prompting (Figure 11), GPT-4o frequently collapses negative and neutral instances into the Neutral class. Few-shot prompting (Figure 12) shifts predictions toward mild and strong positive labels without recovering negative or neutral categories. Figure 9 (c.f. Appendix) illustrates GPT-4o’s tendency to replicate exemplar patterns rather than adapt to the target context, highlighting structural limits of prompt-only evaluation for fine-grained therapeutic assessment.
# In-Context Dim Macro Weighted
F1 Prec. Rec. F1 Prec. Rec.
GPT-4o Zero Shot NJ 23.67 25.81 25.75 23.67 25.81 25.75
W&E 20.44 33.60 24.55 20.44 33.60 24.55
RA 30.99 31.32 34.73 30.99 31.32 34.73
AL 37.20 42.19 33.53 37.20 42.19 33.53
RF 38.94 45.01 37.42 38.94 45.01 37.42
SA 31.68 39.18 30.53 31.68 39.18 30.53
Avg. 30.49 36.18 31.08 30.49 36.18 31.08
Few-Shot NJ 13.01 19.14 25.74 13.01 19.14 25.74
W&E 39.22 44.65 44.91 39.22 44.65 44.91
RA 18.52 45.83 27.24 18.52 45.83 27.24
AL 24.40 38.45 31.13 24.40 38.45 31.13
RF 18.27 27.95 23.65 18.27 27.95 23.65
SA 25.31 33.95 37.42 25.31 33.95 37.42
Avg. 23.12 34.99 31.69 23.12 34.99 31.69
@@ 9 Acknowledgment
Appendix A Baseline Methods
@@
(a) Zero-shot LLMs: GPT-4o Islam and Moushi (2024) and MentalLLaMA Yang et al. (2024) are evaluated without task-specific fine-tuning to gauge the upper bound of out-of-the-box performance. (b) Few-shot LLMs: We further extend this by evaluating GPT-4o under few-shot prompting 8, using carefully constructed in-context examples (Appendix 6) to examine whether performance improves with light in-context supervision. A detailed analysis of both zero-shot and few-shot performance including prompt formats, per-dimension scores, and failure analysis is presented in (Discussion 6). (c) Transformer encoders: MentalBERT Ji et al. (2021), ALBERT Lan et al. (2019), DeBERTa He et al. (2020), and RoBERTa Liu et al. (2019) test how well discriminative, parameter-efficient encoders capture therapeutic nuances when fine-tuned on our data; (d) Encoder–decoders: BART Lewis et al. (2019) and its domain-adapted variant MentalBART Yang et al. (2023) assess the utility of sequence-to-sequence pretraining for dialogue-level classification tasks; (e) Decoder-only LLMs: Qwen3 Yang et al. (2025), LLaMA 3.1/3.2 Grattafiori et al. (2024), Phi-4 Abdin et al. (2025), and Gemma Team et al. (2024) probe the limits of autoregressive generators when repurposed for structured ordinal prediction.
Appendix B Experimental Setup